An info source produces 0 and 1 with probabilities 0.3 and 0.7, respectively. the output of the source is transmitted via a channel that has probabilities of error (turning a 1 into a 0 or a 0 into a 1) equal to 0.2

1) what is the probability that at the output a 1 is observed?

2) What is the probability that a 1 was output of the source if at the output of the channel a 1 is observed.
Answer #1

Answer:

a)

Given,

To determine the probability that at the output a 1 is observed

i.e.,

= probability that 1 is produced and correctly transmitted + probability that 0 is produced and it is transmitted incorrectly

= 0.7 * (1 - 0.2) + 0.3 * 0.2

= 0.56 + 0.06

Required probability = 0.62

b)
Now to determine the probability that 1 was output of the source if at the output of the channel a 1 is observed.

= 0.7*0.8/0.62

= 0.7 * 1.290323

= 0.903225806

= 0.9032